# The Exchangeability Assumption for Permutation Tests of Multiple Regression Models

### v1

@@ 1.1 Exchangeability
In that setting, permuting the treatment variable (taking AP courses or not) creates a null sampling distribution with the unintended side-effect that the treatment variable ($X_{2}$) is no longer associated with income level ($X_{1}$). Because treatment and income are associated in the real data, the permuted data do not retain the observed dependency structure in the original data. That is, the relevant relationship is lost after permutation and the exchangeability condition is violated.
2 Permutation Tests in MLR
2.1 Notation
The population model of interest is
$E[Y]=\beta_{0\cdot 1,2}+\beta_{1\cdot 2}X_{1}+\beta_{2\cdot 1}X_{2},$
where $X_{2}$ is the variable of interest in predicting $Y$, and $X_{1}$ is a nuisance variable. The $\cdot$ (dot) notation indicates the “other” variables included in the model.
The model estimated from the original dataset is given by
$\hat{Y}=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}$
A variable that is permuted is denoted by an asterisk. For example,
$Y^{*}\sim X_{1}+X_{2}$
means that permuted $Y$ values are regressed on the original $X_{1}$ and $X_{2}$. The corresponding model would be based on permuted values; therefore, the estimated coefficients would be indicated by asterisks: $b_{0\cdot 1,2}^{*}$, $b_{1\cdot 2}^{*}$, and $b_{2\cdot 2}^{*}.$
In Section 2.2.3, the permutation happens on residuals from a reduced model. That is, the model is first fit only on $X_{1}$, producing coefficients $b_{0\cdot 1}$ and $b_{1}$. Then residuals are calculated using
$R_{Y\cdot 1}=Y-(b_{0\cdot 1}+b_{1}X_{1}).$
The permuted residuals are added back to the single variable model, arriving at new $Y^{*}$ values which are then regressed on $X_{1}$ and $X_{2},$ the covariates.
$Y^{*}:=b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1}$
2.2 Permuting specifications
Unlike simple linear regression, with MLR there is no obvious choice of how to permute. In what follows, we discuss four different permutation methods and their exchangeability conditions. The goal of permuting is to create a null sampling distribution of the statistic of interest, here $b_{2\cdot 1},$ so thata we may infer the variability of the statistic. That is, a distribution of the statistics under the setting where the response $Y$ and the $X_{2}$ variable are not linearly related. But also, the null distribution needs to be created under the exchangeability condition—that the permuting only leads to a change in the relationship between $Y$ and $X_{2}$ and not to a change to any of the other variable relationships in the linear model. All four of the permuting schemes set the null hypothesis to $H_{0}\!\!:\beta_{2\cdot 1}=0.$ See Manly (1997) for an excellent summary of the different permutation methods within the MLR model.
Statistics that are independent of all unknown parameters are called pivot statistics. In our setting, we use a t-statistic (which is pivotal) as our test statistic for two reasons. First of all, it allows for comparison across all methods, even the method that permutes the residuals of the full model and does not have $\beta_{2\cdot 1}=0$ in the test statistic. Secondly, the advantages of pivot statistics are well-established (Hall and Wilson 1991; Westfall and Young 1993).
2.2.1 Permute $Y$
At first glance, it might seem like permuting the response variable would be a good way to break the relationship between $Y$ and $X_{2}$ (Manly 1986, 1997). Indeed, permuting $Y$ will break the relationship between $Y$ and $X_{2}$ which will force the null hypothesis to be true. However, permuting $Y$ will also break the relationship between $Y$ and $X_{1},$ which may not be acceptable if we need to preserve the relationship to mirror the original data structure. If we permute $Y,$ then exchangeability is met only if $Y$ and $X_{1}$ are uncorrelated in the original dataset.
Summary of permuting $Y$
Regression:
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $Y$. Note that when $Y$ is permuted, the original relationship between $X_{1}$ and $X_{2}$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $Y$, however, is problematic under general conditions where $X_{1}$ and $Y$ are not independent. It is worth pointing out that underlying all of the methods (including Ordinary Least Squares (OLS)), the observations are assumed to be independent of one another.
Table 1: Different permutation schemes, variable relationships that are broken, and variable relationships that are preserved. Note that all methods require the observations to be independent of one another.
Permutation Broken Relationships Preserved Relationships
Permute $Y$ $X_{1}$ $Y$ $X_{1}$ $X_{2}$
Manly (1986), Manly (1997) $X_{2}$ $Y$
Permute $X_{2}$ $X_{1}$ $X_{2}$ $X_{1}$ $Y$
N. Draper and Stoneman (1966) $X_{2}$ $Y$
Permute reduced model residuals $X_{2}$ $Y$ (if $X_{1}$ $X_{2}$ are not highly collinear) $X_{1}$ $X_{2}$
Freedman and Lane (1983) $X_{1}$ $Y$
Permute full model residuals None $X_{1}$ $X_{2}$
ter Braak (1990), ter Braak (1992) $X_{1}$ $Y$
$X_{2}$ $Y$
2.2.2 Permute $X_{2}$
In order to maintain the relationship between $Y$ and $X_{1}$, we might consider permuting $X_{2}$ instead of $Y$ (N. Draper and Stoneman 1966). Indeed, the permutation distribution created from permuting $X_{2}$ will force the null hypothesis to be true. However, permuting $X_{2}$ has the side effect that the relationship between $X_{1}$ and $X_{2}$ will be broken in the permuted data. If the data come from, for example, a randomized clinical trial (where $X_{2}$ is the treatment variable), then $X_{1}$ and $X_{2}$ will be independent in the original dataset, and permuting of $X_{2}$ will not violate the exchangeability condition. If $X_{1}$ and $X_{2}$ are correlated in the original dataset, as in our AP exam and socio-economic status example, then permuting $X_{2}$ violates the exchangeability condition.
Summary of permuting $X_{2}$
Regression:
$Y\sim X_{1}+X_{2}^{*}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting $X_{2}$. Note that when $X_{2}$ is permuted, the original relationship between $X_{1}$ and $Y$ is preserved (which is what we want in terms of exchangeability). The broken relationship between $X_{1}$ and $X_{2}$, however, is problematic under general conditions.
2.2.3 Permute reduced model residuals
In order to address some of the concerns of the first two MLR permutation methods, Freedman and Lane (1983) propose a permutation scheme based on residuals. The permutation preserves the relationship between $X_{1}$ $X_{2}$ as well as the relationship between $X_{1}$ $Y$. However, in order for the relationship between $X_{2}$ $Y$ to be broken (i.e., to obtain a null sampling distribution for the test of $H_{0}:\beta_{2\cdot 1}=0),$ $X_{1}$ and $X_{2}$ must not be highly collinear.
Summary of permuting reduced model residuals
Regression:
$Y^{*}:=b_{0\cdot 1}+b_{1}X_{1}+R^{*}_{Y\cdot 1}$
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b_{2\cdot 1}^{*}-0}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that when permuting the reduced model residuals, the original relationships between both $X_{1}$ and $X_{2}$ and additionally between $X_{1}$ and $Y$ are preserved (which is what we want in terms of exchangeability). The null hypothesis is only true if the relationship between $X_{2}$ and $Y$ is broken, and that happens only when $X_{1}$ and $X_{2}$ are not highly collinear.
Claim. The correlation between $Y^{*}$ and $X_{2}$ is dependent on the correlation between $X_{1}$ and $X_{2}$.
  1. (1)
    The permutation scheme is designed to create new values for the coefficient on $X_{2}$ under the setting that $X_{2}$ and $Y$ are uncorrelated. Repeated sets of the “permuted” $Y^{*}$ values are generated by adding noise (permuted residuals) to the fitted values from the model on $X_{1}$ only. At first glance, it seems as though the method creates $Y^{*}$ values that are associated with $X_{1}$ (because they built from the $X_{1}$-model fitted values) and not associated with $X_{2}$ (because they are built from a model that completely ignores $X_{2}$). However, if $X_{1}$ and $X_{2}$ are highly correlated, then $Y^{*}$ values that are correlated with $X_{1}$ will naturally be correlated with $X_{2}$.
  2. (2)
    In Appendix A we give a proof sketch outlining the derivation for the covariance between $Y^{*}$ and $X_{2}$. Although the proof is only outlined, the dependence on the correlation between $X_{1}$ and $X_{2}$ is clear. The approximate value of $\rho(Y^{*},X_{2})$ derived in Appendix A is validated empirically (results not shown); see Ye (2023) for empirical results.
2.2.4 Permute full model residuals
As an extension to Freedman and Lane (1983), ter Braak’s (1990, 1992) permutation method permutes the residuals from the full model. The method has a bootstrap flavor, but the permutation of the residuals is done without replacement (whereas bootstrapping is done with replacement). The underlying technical condition for ter Braak’s method is that:
$F\bigg{(}\frac{b_{2\cdot 1}^{*}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}\bigg{)}\approx F\bigg{(}\frac{b_{2\cdot 1}-\beta_{2\cdot 1}}{SE(b_{2\cdot 1})}\bigg{)},$ (1)
where $F$ is the cumulative distribution function of the random variable described in the argument of the function.
The sampling distribution of $t^{*}=\frac{b^{*}_{2\cdot 1}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}$ approximates the sampling distribution of $t=\frac{b_{2\cdot 1}-\beta_{2\cdot 1}}{SE(b_{2\cdot 1})}$ due to the theoretical underpinnings from bootstrapping. That is, the variability of $b_{2\cdot 1}^{*}$ around $b_{2\cdot 1}$ mimics the variability of $b_{2\cdot 1}$ around $\beta_{2\cdot 1}.$
Under Equation (1), we can use the $t^{*}$ distribution constructed from many permutations of the same dataset to carry out a hypothesis test. If $H_{0}\!\!:\beta_{2\cdot 1}=0$ is true, then $t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$ would be a likely value in the $t^{*}$ distribution, corresponding to a non-significant p-value, resulting in a failure to reject $H_{0}.$
However, if $H_{A}\!\!:\beta_{2\cdot 1}\neq 0$ is true and, say, $\beta_{2\cdot 1}=47,$ then we would expect $\hat{t}=\frac{b_{2\cdot 1}-47}{SE(b_{2\cdot 1})}$ to lie well within the $t^{*}$ distribution, while $t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$ would lie on the margins, leading to a small p-value that concludes the test by rejecting $H_{0}.$
Furthermore, ter Braak (1990) synthesizes equations from various authors – Efron (1982) for the bootstrap, along with Cox and Hinkley (1974) and Lehmann and D’Abrera (1975) for the permutation – to make the following statements regarding the expected value and variance of the estimated slope coefficients in the bootstrapping versus permutation settings. $b_{2\cdot 1}^{+}$ corresponds to the coefficient derived from a standard bootstrap (full model), while $b_{2\cdot 1}^{*}$ corresponds to the permutation of the residuals under the full model.
$E^{+}(b_{2\cdot 1}^{+})$ $=E^{*}(b_{2\cdot 1}^{*})=b_{2\cdot 1}$ (2)
$var^{+}(b_{2\cdot 1}^{+})$ $=(1-1/n)var^{*}(b_{2\cdot 1}^{*}).$ (3)
Equation (2) indicates that the expected values of $b_{2\cdot 1}^{+}$ and $b_{2\cdot 1}^{*}$ are both $b_{2\cdot 1}$. Equation (3) suggests that the variance of $b_{2\cdot 1}^{+},$ the bootstrapped estimate, is smaller than the variance of $b_{2\cdot 1}^{*}.$ Hence, ter Braak uses Equations (2) and (3), along with the order property that $b_{2\cdot 1}^{+}$ and $b_{2\cdot 1}^{*}$ differ by $O(1/n)$ in second or higher order moments, to justify his proposal of the full model residual permutation strategy.
Summary of permuting full model residuals
Regression:
$Y^{*}:=b_{0\cdot 1,2}+b_{1\cdot 2}X_{1}+b_{2\cdot 1}X_{2}+R^{*}_{Y\cdot 1,2}$
$Y^{*}\sim X_{1}+X_{2}$
Null sampling distribution created using:
$t^{*}=\frac{b^{*}_{2\cdot 1}-b_{2\cdot 1}}{SE(b_{2\cdot 1}^{*})}$
Test statistic:
$t_{obs}=\frac{b_{2\cdot 1}-0}{SE(b_{2\cdot 1})}$
Table LABEL:tbl-perms summarizes the broken and preserved relationships when permuting the reduced model residuals. Note that permuting the residuals under the full model allows all of the exchangeability conditions to hold. The new model does not force the null hypothesis to be true which is why the test statistic measures the deviation of the permuted coefficients (on $X_{2}$) to the original data model coefficient (on $X_{2}$) instead of comparing to zero.
2.3 Why permutation test?
It turns out that even when both the normality conditions and the exchangeability are violated, all four of the aforementioned permutation test methods do reasonably well at controlling Type I errors (Anderson and Legendre 1999; Winkler et al. 2014). Only when the error terms come from extremely skewed distributions do the permutation tests demonstrate substantially higher power than the normal theory methods (Anderson and Legendre 1999). There is strong evidence that in order to violate alpha-level testing, the data must violate the technical conditions to an extreme degree (Anderson and Legendre 1999; Tantawanich 2006; Winkler et al. 2014). Our own simulations (not shown) give the same results that despite differences across exchangeability conditions, the different permutation schemes presented above do not result in substantially different inferential outcomes.
If we anticipate that results will be indistinguishable, why is an understanding of permutation methods (and exchangeability) important?

### v2

@@ 1.2 Motivating example
Breaking the relationship between $Y$ and $X_{2}$ wouldn’t be an issue if $Y$ and $X_{2}$ were not associated. But if $X_{2}$ is a potential confounder of the relationship, then $Y$ and $X_{2}$ may be associated. The permutation procedure where the outcome $Y$ is permuted, therefore, is not exchangeable, even when the null hypothesis of interest is true (i.e., $\beta_{1}=0$). In this setting, permuting the outcome creates a permutation sampling distribution with the unintended consequence that the performance in college ($Y$) is no longer associated with income level ($X_{2}$). Because treatment and income are associated in the real world, the data with permuted $Y$ values do not retain the observed dependency structure that is in the original data. That is, a relevant relationship is lost after a permutation, which ends up violating the exchangeability condition.
Similar issues arise if the predictor $X_{1}$ is permuted (which breaks the relationship between $Y$ and $X_{1}$, as well as between $X_{1}$ and $X_{2}$, but maintains the relationship between $Y$ and $X_{2}$). In either case, more relationships are broken than intended.
What impact does violation of exchangeability have on our inferences? We might consider studies of Type I error rate (when the null hypothesis is true) as well as Type II error rate (when the alternative hypothesis is true). Prior research (Anderson and Legendre 1999; Winkler et al. 2014) has shown that hypothesis testing in MLR is quite robust to the choice of permutation, and, except in cases of extreme error distributions, meeting (or not) the exchangeability conditions (such as breaking the association between $Y$ and $X_{2}$) does not substantially impact the size or power of the test.
However, robustness to exchangeability violations does not hold for all types of exchangeability violations. Consider another example where the underlying data are clustered in some fashion. For example, consider a situation where, for convenience or design purposes, families might be sampled rather than individuals. So instead of having just one individual from each family, we have two students for some families.
One of the conditions for inference in multiple regression is that the observations (or more accurately, the residuals) are independent of one other. In the clustering setting, it may no longer be tenable to assume that individuals within a family are independent of others within the same family.
Principled and flexible methods have been developed for MLR models in the setting where the observations are clustered (see for example Laird and Ware (1982)). Random intercept models, where each observational unit (e.g., family) is assumed to have an underlying (unobserved) random level, are commonly used to account for clustering. However, such mixed effects models are both complicated (mixed effects models are not typically seen in undergraduate curricula) and may require sufficient sample size to assume asymptotic normality of regression coefficients (Maas and Hox 2005).
Permutation methods are particularly desirable with clustered data since they can be adapted to more complicated settings where exchangeability might not be as straightforward as it is in the MLR setting (D. Draper et al. 1993). In the next section we detail how violations of exchangeability in the cluster setting can have a substantial impact on inferential results.
2 When exchangeability affects performance
We revisit our motivating example where we are interested in predicting college performace as the outcome ($Y$) based on a quantitative predictor (family income, $X_{2}$) and a dichotomous treatment (AP course taking, $X_{1}$) where some of the families have two individuals within the sample. An adaption of the permutation tests that accounts for clustering within observational unit is a computational alternative to the mathematical (i.e., based on the Central Limit Theorem) approach, which maintains an appropriate Type I error rate (Winkler et al. 2014, 2015). Here we explore the clustered permutation idea with three different data scenarios, representing independent observations and two different types of clustering that violate the independent observation condition.
2.1 Cluster scenarios
2.1.1 Permuting independent observations
Figure 1 displays an example where the observations are independent and no clustering structure exists. There are eleven subjects, each from a different family; six of the subjects are exposed, $X_{1}$, (e.g., the exposure group takes an AP course) and five of the subjects are control samples (e.g., no AP course). A second variable, $X_{2}$, (e.g., income) is collected on each subject and the outcome variable, $Y$, (e.g., college performance) is measured.
In the example, we permute $X_{1}$, akin to the method of N. Draper and Stoneman (1966). We note that because we are permuting, there are always five exposed observations and six control observations, leading to (${11\choose 6}$ = 462 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a MLR for each permutation use the resulting null distribution of the t-statistics (i.e., the standardized slope coefficients for $X_{1}$) to assess the significance of the treatment variable.
Figure 1: Each color and row represents a family structure. There are eleven individuals, none of whom are in the same family. Note that in the permutation setting, the exposure group and the control group both retain the same number of individuals as in the original family structure.
2.1.2 Permuting homogeneous clustered observations
In the homogeneous clustered (i.e., all individuals within a family have the same value of the treatment variable) data scenario, there are multiple observations of the same type within the observational unit (see Figure 2). There are a total of eight families. Five of the families have a single observation, and three of the families have two individuals.
With homogeneous clustered observations, the permutation scheme must align with the family structure. Here, each family has only one treatment (either exposure or control), so we permute the treatments with the constraint that three of the single-person families and one of the two-person families are control. For the homogeneous clustered example, there are (${5\choose 3}\cdot{3\choose 1}$ = 30 possible permutations (although typically, in larger sample sizes, we won’t render all possible permutations). We fit a MLR for each permutation use the resulting null distribution of the t-statistics (i.e., the standardized slope coefficients for $X_{1}$) to assess the significance of the treatment variable.
Figure 2: Each color and row represents a family structure. There are five single-person “families” and three families with two members each. In the original structure, treatment happens by family. Note that in the permutation setting, the exposure group and the control group both retain the same number of individuals as in the original family structure. The families are permuted together so as to keep intact the original scenario of treatment within family.
2.1.3 Permuting heterogeneous clustered observations
With heterogeneous clustered observations, the permutation scheme must perserve the family structure. Here, the treatment on single-person families will be permuted as if the observations were independent (i.e., a permutation with the constraint that two individuals are exposed and three are control). The permutation of observations in the two-person families happens within the family. That is, with equal probability, the treatment levels are either kept the same or swapped, within each family.
Figure 3: Each color and row represents a family structure. There are five single-person “families” and three families with two members each. In the original structure, for the families of size two, exactly one member of each family is exposed, and the other member of the family is in the control group. Note that in the permutation setting, the exposure group and the control group both end up with the same number of individuals as in the original family structure. For the families of size two, the permutation happens within the family so as to keep intact the original structure of both treatments across the family.
2.2 Cluster simulations
Recall that in the independent MLR setting, even permutation methods that violate exchangeability did not show substantial performance declines (Anderson and Legendre 1999; Winkler et al. 2014). With the clustered family structure, we have constructed a completely different type of exchangeability violation due to the independence (or lack thereof) of observations. The independence exchangeability violation does matter (in terms of performance) as seen in what follows.
We undertook a limited series of simulation studies to explore the behavior of permutation tests when the design structure was intentionally specified to violate the independence condition. The simulations exploited the fact that ignoring clustering for the homogeneous clusters yields standard errors that are too small and that doing so for the heterogeneous cluster design, where family members serve as their own control, yields standard errors that are too large (Cannon et al. 2001).
The heart of our simulations is well described by Winkler et al. (2015), who focus on
For each model/scenario combination, 2000 simulations were generated with $n=20$ subjects per treatment group (more subjects than in Figure 1, Figure 2, and Figure 3). For the clustering scenarios, half of the subjects were singletons and the others shared a family member within the group (homogeneous) or across groups (heterogeneous). The distribution of $X_{1}$ was evenly distributed between 0 and 1. The distribution of $X_{2}$ was specified as normal with mean 0 or 1/2, respectively, with a standard deviation of 1, and an induced correlation between $X_{1}$ and $X_{2}$ of approximately 0.24. The residual standard error was also set to 1. Within each simulation, the null distribution was created using 2000 permutations.
Table LABEL:tbl-simresults displays the results from the simulation study. As expected, the models that do not appropriately account for clustering are anti-conservative (“lm homogeneous”, “naïve permutation homogeneous”) or overly conservative (“lm heterogeneous” and “naïve permutation heterogeneous”); the models are given in boldface. Methods that appropriately account for the clustering (e.g., “lme” or “correct permutation”) include the desired Type I error rate within the 95% CI. When there are cluster dependencies, the naïve method is not advised.
Table 1: Type I error rates for different scenarios (2000 simulations, each with 2000 permutations for each simulation). Models include MLR (lm), multiple linear mixed effects (lme), permutation not accounting for clustering (naïve), and permutation accounting for clustering (correct). Forms of clustering are independent (yellow), homogeneous (white), and heterogeneous (blue). The dichotomous treatment variable was permuted (Draper and Stoneman, 1996). Type I error was computed using $\alpha=0.05$.
Model/Scenario Type I error rate (95% CI) Results
lm independent 0.042 (0.034-0.052) includes desired alpha level
correct permutation independent 0.046 (0.037-0.056) includes desired alpha level
lm homogeneous 0.07 (0.059-0.082) anticonservative
lme homogeneous 0.051 (0.041-0.061) includes desired alpha level
naïve permutation homogeneous 0.072 (0.061-0.084) anticonservative
correct permutation homogeneous 0.045 (0.036-0.055) includes desired alpha level
lm heterogeneous 0.029 (0.023-0.038) conservative
lme heterogeneous 0.044 (0.035-0.054) includes desired alpha level
naïve permutation heterogeneous 0.034 (0.027-0.043) conservative
correct permutation heterogeneous 0.056 (0.047-0.068) includes desired alpha level
3 More ways to carry out Permutation Tests in MLR
The clustering setting demonstrates the importance of understanding exchangeability (and violations). However, we spend the rest of the paper describing exchangeability under the standard independent observation MLR model, because we believe that the discussion that follows is pedagogically quite powerful.
3.1 Why permutation tests?
As previously observed, even when both the normality conditions and the exchangeability are modestly violated, all four of the permutation test methods described below do reasonably well at controlling Type I errors (Anderson and Legendre 1999; Winkler et al. 2014). When error terms come from extremely skewed distributions (e.g., cubed exponentials) the permutation tests maintained the appropriate Type I error rate while normal theory methods were overly conservative (Anderson and Legendre 1999). There is strong evidence that in order to violate alpha-level testing, the data must violate the technical conditions to an extreme degree (Anderson and Legendre 1999; Tantawanich 2006; Winkler et al. 2014). Our own simulations (not shown) give the same results that despite differences across exchangeability conditions, the different permutation schemes (see Section 3.3) do not result in substantially violations of Type I errors.
If we anticipate that results will be indistinguishable, why is an understanding of permutation methods (and exchangeability) important? Why do we present the methods below as important for communicating exchangeability to our students?